# Graph-level Representation Learning with Joint-Embedding Predictive Architectures

### v2

@@ 3.4 Latent Target Prediction
where $cosh(\cdot)$ and $sinh(\cdot)$ are the hyperbolic cosine and sine functions. The predictor module is then tasked with directly locating the target in the unit hyperbola, given the context embedding and the target patch’s positional encoding:
@@
where $W_{n}$ and $b_{n}$ represent the n-th weight matrix and bias vector (i.e., n-th fully connected layer), $\sigma$ is an elementwise non-linear activation function, and $\hat{\psi}^{y}_{l}\in\mathbb{R}^{2}$. This allows us to frame the learning procedure as a regression problem, and the whole network can be trained end-to-end (Fig. 2f). In practice, we use the smooth L1 loss as the distance function, as it is less sensitive to outliers compared to the typical L2 loss (Girshick, 2015):
@@
Thus, we are effectively measuring how far away the context and target patches are in the unit hyperbola of the plane, but the targets are actually described through a high dimensional latent code (Eq. 4). We explicitly show the differences between this choice and using the Euclidean or Hyperbolic distances as energy functions (in the latent space) in Section 4.3. Our proposed pretraining objective forces the context encoder to understand the differences in the hyperbolic angle between the target patches, which can be thought of as establishing an implicit hierarchy between them.
@@
JEPAs are based on a self-distillation procedure. Therefore, they are by definition susceptible to representation collapse (Assran et al., 2023). This is due to the nature of the learning process, where both the context and target representations have to be learned. We formalize this intuition with an example and argue why there is a need to adopt two well-known training tricks that are prevalent in the literature to prevent representation collapse: i) The stop-gradient operation on the target encoder followed by a momentum update (using an Exponential Moving Average (EMA) of the context encoder weights) (Grill et al., 2020; Chen & He, 2021); ii) a simpler parametrization of the predictor compared to the context and target networks (in terms of parameters)(Chen et al., 2020; Baevski et al., 2022);
Let us simplify the problem through the following assumptions: i) The predictor network is linear; ii) There is a one-to-one correspondence between context and target patches. (This holds also in practice due to Eq. 5); iii) The self-predictive task is a least-squares problem in a finite-dimensional vector space. Based on these assumptions, we can rewrite the context features as $X\in\mathbb{R}^{n\times d}$, the target coordinates as $Y\in\mathbb{R}^{n\times 2}$, and the weights of the linear model as $W\in\mathbb{R}^{d\times 2}$. The optimal weights of the predictor are given by solving:
$\operatorname*{arg\,min}\limits_{W}\left\|XW-Y\right\|^{2}$ (7)
@@ 4.1 Experimental setting
We use the TUD datasets (Morris et al., 2020) as commonly done for graph-level SSL (Suresh et al., 2021; Tan et al., 2023). We utilize seven different graph-classification datasets: PROTEINS, MUTAG, DD, REDDIT-BINARY, REDDIT-MULTI-5K, IMDB-BINARY, and IMDB-MULTI. We report the accuracy of ten-fold cross-validation for all classification experiments over five runs (with different seeds). It is worth noting that we retrain the Graph-JEPA model for each fold without ever having access to the testing partition in both the pretraining and fine-tuning stages. We use the ZINC dataset for graph regression and report the Mean Squared Error (MSE) over ten runs (with different seeds), given that the testing partition is already separated. To produce the unique graph-level representations, we feed all the subgraphs through the trained target encoder and then use mean pooling, obtaining a single feature vector $z_{G}\in\mathbb{R}^{d}$ that represents the whole graph. This high-dimensional is then used to fit a linear model with L2 regularization for the downstream task. Specifically, we employ Logistic Regression with L2 regularization on the classification datasets and Ridge Regression for the ZINC dataset. For the datasets that do not natively have node and edge features, we use a simple constant (0) initialization. The subgraph embedding GNN (Figure 2c.) consists of the GIN operator with support for edge features Hu et al. (2019), often referred to as GINE. The neural network modules were trained using the Adam optimizer Kingma & Ba (2014) and implemented using PyTorch Paszke et al. (2019) and PyTorch-Geometric Fey & Lenssen (2019), while the linear classifiers and cross-validation procedure were implemented through the Scikit-Learn library Pedregosa et al. (2011). All experiments were performed on Nvidia RTX 3090 GPUs. Finally, 1 shows the JEPA-specific hyperparameters used for the following experiments.
Table 1: Values of Graph-JEPA specific hyperparameters for the experiments on the TUD datasets.
Hyperparameter PROTEINS MUTAG DD REDDIT-B REDDIT-M5 IMDB-B IMDB-M ZINC
# Subgraphs 32 32 32 128 128 32 32 32
# GNN Layers 2 2 3 2 2 2 2 2
# Encoder Blocks 4 4 4 4 4 4 4 4
Embedding size 512 512 512 512 512 512 512 512
RWSE size 20 15 30 40 40 15 15 20
# context - # target 1 - 2 1 - 3 1 - 4 1 - 4 1 - 4 1- 4 1- 4 1- 4
@@ 4.2 Downstream performance
For the experiments on downstream performance, we follow Suresh et al. (2021) and also report the results of a fully supervised Graph Isomorphism Network (GIN) (Xu et al., 2018), denoted F-GIN in Table 2. We compare Graph-JEPA against four contrastive methods, two generative methods, and one latent self-predictive method (Xie et al., 2022a) (which also regularizes through instance-level reconstruction). As can be seen in Table 2, Graph-JEPA achieves competitive results on all datasets, setting the state-of-the-art as a pretrained backbone on five different datasets and coming second on one (out of eight total). Overall, our proposed framework learns semantic embeddings that work well on different graphs, showing that Graph-JEPA can be utilized as a general pretraining method for graph-level SSL. Notably, Graph-JEPA works well for both classification and regression and performs better than a supervised GIN on all classification datasets. We also provide results with BGRL (Thakoor et al., 2021), a node-level latent self-predictive strategy. We train this model using the official code and hyperparameters and then mean-pool the node representations for the downstream task. The results are underwhelming compared to the models reporting graph-level performance, which is to be expected considering that methods that also perform well on graph-level learning are appropriately designed.
We further explore the performance of our model on the synthetic EXP dataset (Abboud et al., 2020), compared to end-to-end supervised models. This experiment aims to empirically verify if Graph-JEPA can learn highly expressive graph representations (in terms of the commonly used WL hierarchy (Morris et al., 2019)) without relying on supervision. The results in Table 3 show that our model is able to perform much better than commonly used GNNs. Given its local and global exchange of information, this result is to be expected. Most importantly, Graph-JEPA almost matches the flawless performance achieved by He et al. (2023), who train fully supervised.
@@
Graph-JEPA $\underline{75.67}$$\pm$ $\underline{3.78}$ 91.25 $\pm$ 5.75 78.64 $\pm$ 2.35 91.99 $\pm$ 1.59 56.73 $\pm$ 1.96 73.68 $\pm$ 3.24 50.69 $\pm$ 2.91 0.434 $\pm$ 0.014
@@ 4.3 Exploring the Graph-JEPA latent space
As discussed in Section 3.4, the choice of energy function has a big impact on the learned representations. Given the latent prediction task of Graph-JEPA, we expect the latent representations to display hyperbolicity. The predictor network is linearly approximating the behavior of the unit hyperbola such that it best matches the generated target coordinates (Eq. 6). Thus, the network is actually trying to estimate a space that can be considered a particular section of the hyperboloid model (Reynolds, 1993), where hyperbolas appear as geodesics. We are, therefore, evaluating our energy in a restricted part of hyperbolic space. As mentioned before, we find this task to offer great flexibility as it is straightforward to implement and it is computationally efficient compared to the hyperbolic distance used to typically learn hyperbolic embeddings in the Poincaré ball model (Nickel & Kiela, 2017). Table 4 provides empirical evidence for our conjectures regarding the suboptimality of Euclidean or Poincaré embeddings on 4 out of the 8 datasets initially presented in Table 2, where we make sure to choose different graph types for a fair comparison. The results reveal that learning the distance between patches in the 2D unit hyperbola provides a simple way to get the advantages of both embedding types. Hyperbolic embeddings must be learned in lower dimensions due to stability issues (Yu & De Sa, 2021), while Euclidean ones do not properly reflect the dependencies between subgraphs and the hierarchical nature of graph-level concepts. Our results suggest that the hyperbolic (Poincaré) distance is generally a better choice than the Euclidean distance in lower dimensions, but it is computationally unstable and expensive in high dimensions. The proposed approach provides the best overall results. We provide a qualitative example of how the embedding space is altered from our latent prediction objective in Fig. 3.
@@
Table 4: Comparison of Graph-JEPA performance for different distance functions. The optimization for Poincaré embeddings in higher dimensions is problematic, as shown by the NaN loss on the IMDB-B dataset. LD stands for Lower Dimension, where we use a smaller embedding size.
Distance function Ours Euclidean Hyperbolic Euclidean (LD) Hyperbolic (LD)
@@
Table 5: Total training time and model parameters of MVGRL, GraphMAE, and Graph-JEPA for pretraining (single run) based on the optimal configuration for downstream performance. OOM stands for Out-Of-Memory.
@@ 4.4 Additional insights and ablation studies
In an attempt to characterize the efficiency of our proposed model, we perform a simple experimental check. In Table 5, we compare the total training time needed for different Graph-SSL strategies to provide a representation that performs optimally on the downstream task. We show results on the datasets with the largest graphs from Table 2: IMDB and REDDIT-M5. While the runtime is hardware-dependent, all experiments are performed on the same machine. Graph-JEPA displays superior efficiency and promising scaling behavior. The presented runtime is naturally dependent on the self-supervised scheme used in each model, so we do not regard it as a definitive descriptor, but rather an indicator of the potential of fully latent self-predictive models.
@@
Table 6 contains the results of parametrizing the whole architecture, other than the initial GNN encoder, through MLPs. This translates to not using the Attention mechanism at all. For this experiment, and also the following ablations, we consider 4 out of the 8 datasets initially presented in Table 2, making sure to choose different graph domains for a fair comparison. Graph-JEPA still manages to perform well, showing the flexibility of our architecture, even though using the complete Transformer encoders leads to better overall performance and less variance in the predictions.
Table 6: Performance when parametrizing the context and target encoders through MLPs vs using the proposed Transformer encoders.
@@
Following He et al. (2023), it is possible to use the RWSE of the patches as conditioning information. Formally, let $B\in\{0,1\}^{p\times N}$ be the patch assignment matrix, such that $B_{ij}=1$ if $v_{j}\in p_{i}$. We can calculate a coarse patch adjacency matrix $A^{\prime}=BB^{T}\in\mathbb{R}^{p\times p}$, where each $A_{ij}^{\prime}$ contains the node overlap between $p_{i}$ and $p_{j}$. The positional embedding can then be calculated for each patch by simply using the RWSE described in Eq. 1 on $A^{\prime}$. We test Graph-JEPA with these relative positional embeddings and find that they still provide good performance but consistently fall behind the node-level (global) RWSE that we employ in our formulation (Table 7(b)). An issue of these relative patch RWSEs is that the number of shared neighbors can obscure the local peculiarities of each patch, rendering the context given to the predictor more ambiguous.
Random subgraphs.
A natural question that arises in our framework is how to design the spatial partitioning procedure. Using a structured approach like METIS Karypis & Kumar (1998) is intuitive and leads to favorable results. Another option would be to extract random, non-empty subgraphs as context and targets. As seen in Table 7(a), the random patches also provide strong performance, showing that the proposed JEPA architecture is not reliant on the initial input, as is the case for many methods that rely on data augmentation for view generation (Lee et al., 2022). Even though our results show that using a structured way to extract the patches might not be necessary, random sampling can be problematic for larger graphs. Thus, we advocate extracting subgraphs with METIS as it is a safer option in terms of generalizability across different graphs and the inductive bias it provides.
Table 7: (a) Performance when using node-level vs patch-level RWSEs. (b) Performance when extracting subgraphs with METIS vs. using random subgraphs.
@@ 5 Conclusion
In this work, we introduce a new Joint Embedding Predictive Architecture (JEPA) (LeCun, 2022), for graph-level Self-Supervised Learning (SSL). An appropriate design of the model, both in terms of data preparation and pretraining objective, reveals that it is possible for a neural network to self-organize the semantic knowledge embedded in a graph, demonstrating competitive performance in different graph data and tasks. Future research directions include extending the proposed method to nodes and edge-level learning, theoretically exploring the expressiveness of Graph-JEPA, and gaining more insights into the optimal geometry of the embedding space for graph SSL.

### v3

@@ 3.4 Latent Target Prediction
where $cosh(\cdot)$ and $sinh(\cdot)$ are the hyperbolic cosine and sine functions respectively. The predictor module is then tasked with directly locating the target in the unit hyperbola, given the context embedding and the target patch’s positional encoding Fig 2e):
@@
where $W_{n}$ and $b_{n}$ represent the n-th weight matrix and bias vector (i.e., n-th fully connected layer), $\sigma$ is an elementwise non-linear activation function, and $\hat{\psi}^{y}_{l}\in\mathbb{R}^{2}$. This allows us to frame the learning procedure as a low-dimensional regression problem, and the whole network can be trained end-to-end (Fig. 2f). In practice, we use the smooth L1 loss as the distance function, as it is less sensitive to outliers compared to the typical L2 loss (Girshick, 2015):
@@
Thus, we are effectively measuring how far away the context and target patches are in the unit hyperbola of the plane by first describing the subgraphs through a high dimensional latent code (Eq. 4). This representation balances the expressivity of high-dimensional embeddings and the hierarchical organization provided by hyperbolic geometry. Intuitively, the hyperbolic angle $\alpha$ represents an aggregate measure of the subgraph’s latent code, compressed into a scalar value. In high dimensions, this scalar value will concentrate around one value, which provides the base level of the hierarchy. At the same time, the encoder networks have the freedom to produce largely deviant latent codes that lead to higher hierarchy levels. A visualization of this process is provided in the Appendix (Fig. 4). We explicitly show the differences between this choice and using the Euclidean or (Poincaré) Hyperbolic distance as energy functions for the training procedure in Sec. 4.3. Our proposed pretraining objective forces the context encoder to understand the differences in the hyperbolic angle between the target patches, which can be thought of as establishing an implicit hierarchy between them.
@@
JEPAs are based on a self-distillation procedure. Therefore, they are by definition susceptible to representation collapse (Assran et al., 2023). This is due to the nature of the learning process, where both the context and target representations have to be learned. We formalize this intuition with an example and argue why there is a need to adopt two well-known training tricks that are prevalent in the literature to prevent representation collapse: i) The stop-gradient operation on the target encoder followed by a momentum update (using an Exponential Moving Average (EMA) of the context encoder weights) (Grill et al., 2020; Chen & He, 2021); ii) a simpler parametrization of the predictor compared to the context and target networks (in terms of parameters)(Chen et al., 2020; Baevski et al., 2022). Let us simplify the problem through the following assumptions: i) The predictor network is linear; ii) There is a one-to-one correspondence between context and target patches (this holds in practice due to Eq. 5); iii) The self-predictive task is a least-squares problem in a finite-dimensional vector space. Based on these assumptions, we can rewrite the context features as $X\in\mathbb{R}^{n\times d}$, the target coordinates as $Y\in\mathbb{R}^{n\times 2}$, and the weights of the linear model as $W\in\mathbb{R}^{d\times 2}$. The optimal weights of the predictor are given by solving:
$\operatorname*{arg\,min}\limits_{W}\left\|XW-Y\right\|^{2},$ (7)
@@ 4.1 Experimental Setting
We use the TUD datasets (Morris et al., 2020) as commonly done for graph-level SSL (Suresh et al., 2021; Tan et al., 2023). We utilize seven different graph-classification datasets: PROTEINS, MUTAG, DD, REDDIT-BINARY, REDDIT-MULTI-5K, IMDB-BINARY, and IMDB-MULTI. We report the accuracy of ten-fold cross-validation for all classification experiments over five runs (with different seeds). It is worth noting that we retrain the Graph-JEPA model for each fold without ever having access to the testing partition in both the pretraining and fine-tuning stages. We use the ZINC dataset for graph regression and report the Mean Squared Error (MSE) over ten runs (with different seeds), given that the testing partition is already separated. To produce the unique graph-level representations, we feed all the subgraphs through the trained target encoder and then use mean pooling, obtaining a single feature vector $z_{G}\in\mathbb{R}^{d}$ that represents the whole graph. This vector representation is then used to fit a linear model with L2 regularization for the downstream task. Specifically, we employ Logistic Regression with L2 regularization on the classification datasets and Ridge Regression for the ZINC dataset. For the datasets that do not natively have node and edge features, we use a simple constant (0) initialization. The subgraph embedding GNN (Fig. 2c.) consists of the GIN operator with support for edge features (Hu et al., 2019), often referred to as GINE. The neural network modules were trained using the Adam optimizer (Kingma & Ba, 2014) and implemented using PyTorch (Paszke et al., 2019) and PyTorch-Geometric (Fey & Lenssen, 2019), while the linear classifiers and cross-validation procedure were implemented through the Scikit-Learn library (Pedregosa et al., 2011). All experiments were performed on Nvidia RTX 3090 GPUs. Tables 7 and 8 in the Appendix contain the dataset statistics and the JEPA-specific hyperparameters used in the following experiments, respectively.
@@ 4.2 Downstream Performance
For the experiments on downstream performance, we follow Suresh et al. (2021) and also report the results of a fully supervised Graph Isomorphism Network (GIN) (Xu et al., 2018), denoted F-GIN in Tab. 1. We compare Graph-JEPA against four contrastive methods, two generative methods, and one latent self-predictive method (Xie et al., 2022a) (which also regularizes through instance-level reconstruction). As can be seen in Tab. 1, Graph-JEPA achieves competitive results on all datasets, setting the state-of-the-art as a pretrained backbone on five different datasets and coming second on one (out of eight total). Overall, our proposed framework learns semantic embeddings that work well on different graphs, showing that Graph-JEPA can be utilized as a general pretraining method for graph-level SSL. Notably, Graph-JEPA works well for both classification and regression and performs better than a supervised GIN on all classification datasets. We also provide results with BGRL (Thakoor et al., 2021), a node-level latent self-predictive strategy. We train this model using the official code and hyperparameters and then mean-pool the node representations for the downstream task. The results are underwhelming compared to the models reporting graph-level performance, which is to be expected considering that methods that also perform well on graph-level learning are appropriately designed.
One notable aspect of the downstream performance of Graph-JEPA is the variance in the results. We conjecture that this variance in the results is due to three main factors. First, the randomness inherent in the subgraph partitioning process can lead to the selection of trivial context subgraphs, making it difficult to accurately predict latent representations of the target subgraphs. Second, while the JEPA paradigm has proven effective in the graph domain, it is susceptible to representation collapse, mainly when the previously mentioned issue is combined with positional embeddings of smaller, similar subgraphs. This would lead to nearly identical latent representations and, thus, reduced downstream performance. The issue is also compounded due to the fact that we did not extensively tune hyperparameters, as our focus was on demonstrating the architecture’s efficacy with a simple training regimen. Lastly, in contrast to standard protocols that involve training on the entire dataset before fine-tuning, we trained the self-supervised model exclusively on the training set to ensure fairness. Despite this, average-case performance remains the most crucial aspect of our experimental validation.
We further explore the performance of our model on the synthetic EXP dataset (Abboud et al., 2020), compared to end-to-end supervised models. This experiment aims to empirically verify if Graph-JEPA can learn highly expressive graph representations (in terms of the commonly used WL hierarchy (Morris et al., 2019)) without relying on supervision. The results in Tab. 2 show that our model is able to perform much better than commonly used GNNs. Given its local and global exchange of information, this result is to be expected. Most importantly, Graph-JEPA almost matches the flawless performance achieved by He et al. (2023), who train fully supervised.
@@
Graph-JEPA $\underline{75.68}$$\pm$ $\underline{3.78}$ 91.25 $\pm$ 5.75 78.64 $\pm$ 2.35 91.99 $\pm$ 1.59 56.73 $\pm$ 1.96 73.68 $\pm$ 3.24 50.69 $\pm$ 2.91 0.434 $\pm$ 0.014
@@ 4.3 Exploring the Graph-JEPA Latent Space
As discussed in Sec. 3.4, the choice of energy function has a big impact on the learned representations. Given the latent prediction task of Graph-JEPA, we expect the latent representations to display hyperbolicity. The predictor network is linearly approximating the behavior of the unit hyperbola such that it best matches the generated target coordinates (Eq. 6). Thus, the network is actually trying to estimate a space that can be considered a particular section of the hyperboloid model (Reynolds, 1993), where hyperbolas appear as geodesics. We are, therefore, evaluating our energy in a restricted part of hyperbolic space. As mentioned before, we find this task to offer great flexibility as it is straightforward to implement and it is computationally efficient compared to the hyperbolic distance used to typically learn hyperbolic embeddings in the Poincaré ball model (Nickel & Kiela, 2017). Tab. 3 provides empirical evidence for our conjectures regarding the suboptimality of Euclidean or Poincaré embeddings on 4 out of the 8 datasets initially presented in Tab. 1, where we make sure to choose different graph types for a fair comparison. The results reveal that learning the distance between patches in the 2D unit hyperbola provides a simple way to get the advantages of both embedding types. Hyperbolic embeddings must be learned in lower dimensions due to stability issues (Yu & De Sa, 2021), while Euclidean ones do not properly reflect the dependencies between subgraphs and the hierarchical nature of graph-level concepts. Our results suggest that the hyperbolic (Poincaré) distance is generally a better choice than the Euclidean distance in lower dimensions, but it is computationally unstable and expensive in high dimensions. The proposed approach provides the best overall results. We provide a qualitative example of how the embedding space is altered from our latent prediction objective in Fig. 3.
@@
Table 3: Comparison of Graph-JEPA performance for different distance functions. The optimization for Poincaré embeddings in higher dimensions is problematic, as shown by the NaN loss on the IMDB-B dataset. LD stands for Lower Dimension, where we use a smaller embedding size. The best result is reported in boldface.
Dataset Ours Euclidean Hyperbolic Euclidean (LD) Hyperbolic (LD)
@@
Table 4: Total training time and model parameters of MVGRL, GraphMAE, and Graph-JEPA for pretraining (single run) based on the optimal configuration for downstream performance. OOM stands for Out-Of-Memory. The best result is reported in boldface.
@@ 4.4 Additional Insights and Ablation Studies
In an attempt to characterize the efficiency of our proposed model, we perform a simple experimental check. In Tab. 4, we compare the total training time needed for different Graph-SSL strategies to provide a representation that performs optimally on the downstream task. We show results on the datasets with the largest graphs from Tab. 1: IMDB and REDDIT-M5. Since runtime is hardware-dependent, all experiments are performed on the same machine. Graph-JEPA displays superior efficiency and promising scaling behavior. The presented runtime is naturally dependent on the self-supervised scheme used in each model, so we do not regard it as a definitive descriptor but rather an indicator of the potential of entirely latent self-predictive models.
@@
Tab. 5 contains the results of parametrizing the whole architecture, other than the initial GNN encoder, through Multilayer Perceptrons (MLPs). This translates to not using the Attention mechanism at all. For this experiment, and also the following ablations, we consider 4 out of the 8 datasets initially presented in Tab. 1, making sure to choose different graph domains for a fair comparison. Graph-JEPA still manages to perform well, showing the flexibility of our architecture, even though using the complete Transformer encoders leads to better overall performance and less variance in the predictions.
Table 5: Performance when parametrizing the context and target encoders through MLPs vs using the proposed Transformer encoders. The best result is reported in boldface.
@@
Following He et al. (2023), it is possible to use the RWSE of the patches as conditioning information. Formally, let $B\in\{0,1\}^{p\times N}$ be the patch incidence matrix, such that $B_{ij}=1$ if $v_{j}\in p_{i}$. We can calculate a coarse patch adjacency matrix $A^{\prime}=BB^{T}\in\mathbb{R}^{p\times p}$, where each $A_{ij}^{\prime}$ contains the node overlap between $p_{i}$ and $p_{j}$. The positional embedding can be calculated for each patch using the RWSE described in Eq. 1 on $A^{\prime}$. We call such an embedding relative, as it can only capture positional differences in a relative manner between patches. We test Graph-JEPA with these relative positional embeddings and find that they still provide good performance but consistently fall behind the node-level (global) RWSE that we employ in our formulation (Tab. 6(b)). An issue of these relative patch RWSEs is that the number of shared neighbors can obscure the local peculiarities of each patch, rendering the context given to the predictor more ambiguous.
Random Partitioning.
A natural question in our framework is how to design the spatial partitioning procedure. Using a structured approach like METIS (Karypis & Kumar, 1998) is intuitive and leads to favorable results. Another option would be to extract random, non-empty subgraphs as context and targets. As seen in Tab. 6(a), the random patches also provide strong performance, showing that the proposed JEPA architecture is not reliant on the initial input, which is the case for many methods that rely on data augmentation for view generation (Lee et al., 2022). Even though our results show that using a structured way to extract the patches might not be necessary, it is an idea that generalizes well to different graph types and sizes. Thus, we advocate extracting subgraphs with METIS as it is a safer option in terms of generalizability across different graphs and the inductive bias it provides.
Table 6: (a) Performance when extracting subgraphs with METIS vs. using random subgraphs. (b) Performance when using node-level vs patch-level RWSEs. The best result is reported in boldface.
@@ 5 Conclusion
In this work, we introduce a new Joint Embedding Predictive Architecture (JEPA) for graph-level Self-Supervised Learning (SSL). An appropriate design of the model, both in terms of data preparation and pretraining objective, reveals that it is possible for a neural network to self-organize the semantic knowledge embedded in a graph, demonstrating competitive performance in different graph data and tasks. Future research directions include extending the proposed method to node and edge-level learning, theoretically exploring the expressiveness of Graph-JEPA, and gaining more insights into the optimal geometry of the latent space for general graph SSL.
Acknowledgments
Geri Skenderi is funded by the European Union through the Next Generation EU - MIUR PRIN PNRR 2022 Grant P20229PBZR. The views and opinions expressed are, however, those of the authors only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible.